Please help w/ the following: Thank you

The amount of money A accured at the end of N years when a certain amount P is invested at a compound annual rate r is given by A=P(1+r)^n. If a person invests $160 at 4% interest compounded annually, find the approximate amount obtained at the end of 15 years.

Answer by checkley71(8403)

A=160(1+.04)^15 OR A=160(1.04)^15 OR A=160(1.8) OR A=$288